Use the Desmos kept grabbing calculator to generate the least squares regression line and statistics for the data in the table. What is the residual value for X =18

X Y

2 3

4 9

11 12

15 17

18 16

25 22


Answers:

-0.4216

-1.2389

0.4216

2.5784


1
Expert's answer
2021-03-15T08:12:49-0400
"y=mx+b"

"m=0.7404, b=3.9117"


"y=0.7404x+3.9117"

"x=18"


"y=0.7404(18)+3.9117=17.2389"

Residual


"16-17.2389= -1.2389"

-1.2389